OpenStudy (anonymous):

Find the hypotenuse of a right triangle with legs of 9 inches and 12 inches

OpenStudy (anonymous):

You would use the formula a^2 + b^2 = c^2 9^2 + 12^2 = c^2 81 + 144 = c^2 225 = c^2 c = 15 The hypotenuse is 15 inches